I am 48, my brother is 42. What is the equation which models how old, in years, I will be when I am 3 times as old as my younger brother.
Answer by stanbon(75887)

I am 48, my brother is 42. What is the equation which models how old, in years, I will be when I am 3 times as old as my younger brother.
==============
Let the missing years be "x".
Equation:
(48+x) = 3(42+x)
48+x = 126 + 3x
2x = -78
x = -39
-----------------
Your age then: 48-39 = 9
Your brother's age then: 42-39 = 3
-------------------------------------------
Cheers,
Stan H.